OpenStudy (ashy98):

What is the sum of the arithmetic sequence 3, 9, 15..., if there are 26 terms? A.) 2,028 B.) 1,452 C.) 2,268 D.) 1,728

OpenStudy (mathstudent55):

1. find the 26th term 2. add the first 26 terms

OpenStudy (mathstudent55):

What is the common difference?

OpenStudy (mathstudent55):

To find the common difference, subtract a term from the next term.

OpenStudy (mathstudent55):

d = common difference = 9 - 3 = 15 - 9 What is d?

OpenStudy (ashy98):

the answer would be A

OpenStudy (mathstudent55):

Formula to find the nth term: \(\large a_n = a_1 + d(n - 1)\) Formula to find the sum of the first n terms: \(\large S_n = \dfrac{n(a_1 + a_n)}{2} \) Using the formulas above, you get: \(a_n = 153\) \(S_n = 2028\)

OpenStudy (mathstudent55):

You are correct.

OpenStudy (ashy98):

okay thank you!(: